the average age of my father, mother, aunt, cousin, niece, uncle, sister and myself is 25 years what was the total 3 years ago?
Found 2 solutions by ankor@dixie-net.com, MathTherapy:
Answer by ankor@dixie-net.com(22740)

the average age of my father, mother, aunt, cousin, niece, uncle, sister and myself is 25 years
what was the total 3 years ago?
:
The total years of 8 people with an average age of 25:
8 + 25 = 200
the reduction of the age of of 8 people 3 yrs ago
8 * 3 = 24
therefore
200 - 24 = 176 yrs

Answer by MathTherapy(10552)


the average age of my father, mother, aunt, cousin, niece, uncle, sister and myself is 25 years what was the total 3 years ago?
25(8) - 3(8) = 22(8) = 176